OpenStudy (anonymous):

solve sinθ+sin3θ=sin2θ for 0<θ<180

mathslover (mathslover):

Hello @jimswig88 . Let me first make sure that do you know the formula for \(\sin 3 \theta \) and \(\sin 2 \theta\)

OpenStudy (anonymous):

Is sin3θ =sin2θcosθ+cos2θsin? or is that a mistake

mathslover (mathslover):

wait! I think I have to re-login here

mathslover (mathslover):

I am back .

OpenStudy (mertsj):

\[\sin 3\theta =\sin (2\theta+\theta)=\sin 2\theta \cos \theta+\cos 2\theta \sin \theta\]

mathslover (mathslover):

Write : \(\large{\sin 3 \theta = \sin (2\theta + \theta )}\)

mathslover (mathslover):

And then use : \(\large{\color{blue}{\sin(A+B)}}\) . I hope you know the formula for sin(A+B) ?

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

mathslover (mathslover):

@jimswig88 can u tell me the formula for : \(\large{\color{red}{\sin(A+B)}}\)

OpenStudy (anonymous):

sinAcosB+sinBcosA?

mathslover (mathslover):

Right . So what will be : \(\large{\color{grey}{\sin(2\theta + \theta)}}\) ?

OpenStudy (anonymous):

Unsure were to go from hre

mathslover (mathslover):

See : \(\large{\color{blue}{ \sin(2\theta +\theta)} = \color{red}{\sin 2\theta \cos \theta + \cos 2\theta \sin \theta }}\) right?

OpenStudy (anonymous):

Oh ok !makes sense

mathslover (mathslover):

Now , we have : \(\large{\sin \theta + \sin 3\theta = \sin 2\theta}\) What if i put the formula up there for sin 2 theta also

OpenStudy (anonymous):

Double angle formula yes?

mathslover (mathslover):

yes. Let us both try to do that .

OpenStudy (anonymous):

2sinθcosθ?

mathslover (mathslover):

yep

OpenStudy (anonymous):

oh good stuff. So i have sinθ+sin2θcosθ+cos2θsinθ=2sinθcosθ. How is that looking?

mathslover (mathslover):

Oh that looks scary :P

OpenStudy (anonymous):

Ha yes it does

mathslover (mathslover):

I got it : we have \(\large{\sin \theta + \sin 3 \theta = \sin 2\theta}\) why dont' use sin A + sin B formula ?

mathslover (mathslover):

Do you know the formula for sin A + sin B ?

OpenStudy (anonymous):

I can not find it in my notes. I know its simple arrrghhh!

mathslover (mathslover):

\[\large{\color{blue}{\sin A + \sin B } = \color{orange}{ 2\sin (\frac{A+B}{2}) \cos (\frac{A-B}{2})}}\]

mathslover (mathslover):

Now let's put : A = 2 theta and B = theta.

OpenStudy (anonymous):

Oh yes thats the one! not the one i was thinking then ha

OpenStudy (anonymous):

sum to product formula yes remember now

mathslover (mathslover):

I made the mistake in previous one . Sorry for that.

OpenStudy (anonymous):

its ok any help is great thanks

mathslover (mathslover):

Now, \[\large{\textbf{we have }}\] : \[\large{2\sin (2 \theta) \cos \theta = \sin 2\theta}\] Right?

OpenStudy (anonymous):

The 2sin(2θ)cosθ being equivalent to sin3θ or all of it. as in being equivalent to sinθ+sin3θ?

mathslover (mathslover):

^ that is equivalent to \(\large{\sin \theta + \sin 3\theta}\)

OpenStudy (anonymous):

yes got you

mathslover (mathslover):

\[\large{2\sin 2\theta \cos \theta = \sin \theta + \sin 3\theta}\] and \(\large{\sin \theta + \sin 3\theta = \sin 2\theta}\) thus : \(\large{\color{blue}{2\sin 2\theta \cos \theta = \sin 2\theta}}\)

mathslover (mathslover):

Now what do you think should be the next step?

OpenStudy (anonymous):

change sin2θ into double angle 2sinθcosθ?

mathslover (mathslover):

No. See can you divide sin 2 theta both sides?

OpenStudy (anonymous):

oh ok so would that leave 2cosθ=1

mathslover (mathslover):

Yes u can get that (but note if theta can be equal to zero then you can not divide sin 2 theta both sides as 0/0 is indeterminate) . therefore \(\large{\cos \theta= \frac{1}{2}}\) and also : \(\large{2\sin 2\theta \cos \theta - \sin 2\theta = 0}\) \(\large{\color{blue}{2\sin 2\theta} (\color{red}{\cos \theta - 1} ) = 0}\) \(\large{\color{orange}{2\sin 2\theta = 0}}\) thus sin 2 theta = 9 or cos theta = 1/2

mathslover (mathslover):

or \(\large{\sin 2\theta (2\cos \theta -1) = 0}\) => \(\sin 2\theta = 0\) or \(\cos \theta = \frac{1}{2}\) got it?

OpenStudy (anonymous):

that is fantastic thank you very much. the step by step guide has been great

mathslover (mathslover):

So did you like : Open + Study ?

OpenStudy (anonymous):

yes first time on here and been a big help

mathslover (mathslover):

Great! So I hope that you will keep asking your doubts here and also do not forget to help others ... :) \[\huge{\color{orange}{\mathbb{WELCOME}} \space \color{red}{\textbf{TO}} \space \frak{\color{red}{OPENSTUDY}}}\]

OpenStudy (anonymous):

Thanks again i will make sure i will help the level 2 maths. Lot easier

mathslover (mathslover):

You're welcome and sure that will be great.